OpenStudy (anonymous):

Which of the following is the solution set of 18 - 3n + 2 = n + 20 - 4n? Ø {0} all reals

OpenStudy (anonymous):

18 - 3n + 2 = n + 20 - 4n 20-3n=20-3n 3n=3n hence n belongs to all reals